The measured pH of the H2SO4 solution in water is 1.8. Calculate the volume of water needed to add to 1.00 L of this solution to increase the pH to 2.55 (The pH scale problem)




1
Expert's answer
2022-05-12T02:30:04-0400

"H_2SO_4 \rightarrow 2H^++SO_4^-"

pH=-ln([H+])

2.55=-ln([H+])

[H+]=10-2.55=0.00282 M - we need

1.8=-ln([H+])

[H+]=10-1.8=0.01585 M - we have

"0.00282=0.01585\/V_{solution}"

"V_{solution}=0.01585\/0.00282=5.62 l"

"V_{solution}=V_{initial} + V_{add}"

"V_{add}=V_{solutuon}-V_{inittlial}=5.62-1=4.62 l"

Answer: We need to add 4.62 l of water.